Projection Admin 10/23

To do this evening/ was completed: A lot of teching and a lot of meetin' and greetin'! Whew. See all notes below, but for tonight (when there's time!) please start building more SE playlists. That would be enormously helpful! Thank you! Can't wait to get rid of some more stuff on the server tomorrow... - Lillian

Presentation Issues:

Preshow Clips/Playlist: Ran through all the new playlists and their content - Frankenstein (I slightly adjusted volume in one preshow segment for this), Springsteen all techd, Shelby Oaks and Blue Moon sorted as well. Now, I did notice that the A1 playlists had the scope version of the trailer for Sentimental Value (please be careful when ingesting/building!) so I switched it to Flat. - L

Shleby Oaks feature should be in Scope. Playlists updated.  L

Frankenstein feature is not Scope, and now is in Flat. Updated. - L

Trailer Updates:

Special Events/ AV set up: Finished the setup for KGO tonight - mics are charging. Added a new spots setting for Stage Right so there's less bleed on the screen (looks nicer this way for the presentation. Kevin is pleased). I already went through the ROS with Will and printed out a tech version I edited with all the cues he'll need (spots/lights/volumes, etc). Kevin and Camila and I tech'd this with Cristina earlier today. The preshow is all good to go (playlist is scheduled and it'll be an easy changeover to the laptop from there). Mic stands sorted, etc. L

Setup the microphone and tech'd the playlist for Return of the Living Dead private event in T6 tonight. - L

Transferred the SAGG shorts to T3 this morning. Cristina and I went through more than half of the shorts, but we still need to look at a few more. I have my notes. Everything's looking just fine. Last short we needed has been ingested as per this evening. - L

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): DCP'd and ingested Worm shorts for SAGG. - L

Outgoing/Shipping (content): Warriors out. I'll take care of Scooby tomorrow. - L

Print/video Inspection: Ran Innocent Blood for Cristina and gosh does it look great! Volume noted. - L

Asset tracking doc updated/checked?:

Maintenance/cleaning: Alec cleaned the platter top to bottom and says feel free to practice as much as you have time for! The Oppenheimer notes are still on the wall and are a great reference point (if I do say so myself) but there's additional notes on the drive. Now, if you don't feel comfortable doing that on your own, we will absolutely spend time practicing together during admin shifts! - L

Cue sheets updated/checked (every shift): Adjusted! Please be sure to double check everything before you put your initials down. 

Marquee updated?: We'll do a big haul tomorrow as today is the last day for many titles. 

Posters updated?: Alec adjusted the posters. 

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will